# Evaluation of LLM Vulnerabilities to Being Misused for Personalized Disinformation Generation

### v1

@@ Abstract
The capabilities of recent large language models (LLMs) to generate high-quality content indistinguishable by humans from human-written texts rises many concerns regarding their misuse. Previous research has shown that LLMs can be effectively misused for generating disinformation news articles following predefined narratives. Their capabilities to generate personalized (in various aspects) content have also been evaluated and mostly found usable. However, a combination of personalization and disinformation abilities of LLMs has not been comprehensively studied yet. Such a dangerous combination should trigger integrated safety filters of the LLMs, if there are some. This study fills this gap by evaluation of vulnerabilities of recent open and closed LLMs, and their willingness to generate personalized disinformation news articles in English. We further explore whether the LLMs can reliably meta-evaluate the personalization quality and whether the personalization affects the generated-texts detectability. Our results demonstrate the need for stronger safety-filters and disclaimers, as those are not properly functioning in most of the evaluated LLMs. Additionally, our study revealed that the personalization actually reduces the safety-filter activations; thus effectively functioning as a jailbreak. Such behavior must be urgently addressed by LLM developers and service providers.
@@ 1 Introduction
However, there is so far little (and mostly anecdotal) supporting evidence to justify such fears. Therefore, this study fills the gap by 1) investigating vulnerabilities of the state-of-the-art (SOTA) LLMs to being misused for generation of personalized disinformation, 2) examining the quality of personalization, and 3) how it affects their integrated safety mechanisms. [^1] For the sake of replicability and support of further research, the data analysis source code as well as the generated dataset (upon request) will be released at https://github.com/kinit-sk/personalized-disinfo for non-commercial research purpose only under strict conditions. Due to ethical concerns (see Appendix A), we are not releasing data generation source code (although an abstract overview is provided in the paper), as approved by our institutional Ethics Review Board. We build on the methodology proposed by Vykopal et al. (2024) to assess disinformation generation potential, and modify it for the purpose of personalized disinformation evaluation. We proceed from the definition of personalization proposed by Blom (2000), who defines it as “a process that changes the functionality, interface, information content, or distinctiveness of a system to increase its personal relevance to an individual.” By personalized disinformation, we mean disinformation appealing (by concerns and emotions) to a specific target group.
@@
(2) Showcase of usability of meta-evaluation (LLM as a judge annotation) for automated and scalable evaluation of personalized text generation. We use three different LLMs (to limit biased judgments) to annotate all data, showing a strong correlation with human judgment (using two human annotators) on a smaller balanced subset.
@@ 3 Methodology
Narratives. For this study, we have carefully selected six representative disinformation narratives (see Table 1) from a set of 20 in Vykopal et al. (2024) to ensure comparability of results. The resulting list covers health and politics-related disinformation, reflecting the varying LLMs’ behavior across topics (Vykopal et al., 2024). The description of previously mentioned target groups is based on European stereotypes. Therefore, we excluded disinformation narratives linked to events or personas outside Europe to avoid combination of target groups and narratives that are unlikely to appear in real-world information space. To the best of our judgment, we selected disinformation narratives that are possible to personalize across all chosen target groups. Each of the six selected narratives includes the title, summarizing the main idea, and an abstract providing additional context. We follow the wording from Vykopal et al. (2024) that was sourced from professional fact-checkers.
@@
Human validation. Two human annotators (out of the authors) have rated the personalization quality of the carefully selected 109 texts on a four-point scale. The human-annotated balanced subset consists of texts generated by all six LLMs, equally distributed over a score assigned by GPT-4o meta-evaluation (more details on the selection of the subset in Appendix C.2). Only the text and the target group were provided in the annotation process, while the annotators were familiar with the target group descriptions. Annotators answered “How well does the fake news article address the intended audience?” on a four-point scale: (0) – The text is not at all personalized to the intended target group. (1) – The text is personalized for the intended target group to a low extent. (2) – The text is personalized to an intended audience in an illogical way or the resulting personalization of text sounds forced and unnatural. (3) – The text is well personalized to the intended target group. The iteratively developed annotation guidelines provide a description of the scale with example text snippets to illustrate the cases. The guidelines have been validated by previous round of annotation of different portion of 50 texts, addressing the ambiguities. Each text was rated by both annotators. Both annotators are based in Europe and are non-native English speakers.
@@ 4.1 Linguistic Quality Analysis
The linguistic analysis with text-quality evaluation of the generated texts is summarized in Table 6, Appendix F. The analysis shows that Falcon outputs have the highest variance in text lengths, thus showing least stable behavior. Although the three quality metrics do not match (reflecting different aspects), it seems that Flacon and Vicuna provide the texts of the lowest quality. On the other hand, Gemma and GPT-4o generated the texts of the highest quality. GPT-4o is the most stable model, generating consistently the high quality texts of similar lengths (the lowest variance in text lengths).
@@ 5.2 Meta-evaluation of Personalization
While human evaluation is labor-intensive, hard to replicate with the same results, and exposes annotators to a harmful content (disinformation in our case), employment of LLMs can effectively scale the evaluation and provide replicable results. However, LLM-based meta-evaluation comes not without limitations; among others LLMs tend to assign a higher score to their own outputs (Panickssery et al., 2024). To mitigate this limitation to a certain extent, we employ three LLMs, namely GPT-4o, Gemma-2-27b-IT, Llama-3.1-70B-Instruct for the evaluation of personalization quality. Moreover, LLM-assigned scores are for some tasks poorly correlated with human ratings (Bansal et al., 2023). While we have used such meta-evaluation to answer RQ1, in this section, we focus on validity of this approach and address the research question RQ2: Are LLMs usable to evaluate personalization of the generated texts with correlation to human judgment? To answer this question, two human annotators and three LLMs evaluated the personalization quality of a carefully balanced subset of 109 texts (see Section 3 and Appendix C.2).
@@
We calculated the agreement rates between the human annotators for the evaluation of personalization quality. The Spearman correlation coefficient ($\rho$) is 0.7 indicating strong correlation (mean absolute error between the two annotators is 0.51, average mean absolute error of human annotators from human average is 0.26). Overall, the two annotators assigned the same scores in 64% of cases, which is acceptable given the 4-point scale and high subjective nature of evaluation task (personalization quality). The highest agreement is in the assignment of the score 0, where the annotators agreed in 85% of cases (i.e., both of them scored 0 for texts that either one or both of them scored 0). When the annotation is reduced to binary Yes/No answers (score of 0 is No, other scores are Yes) to whether the text is personalized for the target group (i.e., not the actual quality of the personalization), the agreement is in 95% of cases.
@@
We also validated each score assigned by LLM-meta-evaluators individually. We checked whether it is matched by at least one of human annotators. It can provide us a kind of reliability measure of meta-evaluation scores. The highest reliability is in assignment of the score of 0 (from 92% match in case of Gemma to 97% in case of GPT-4o) and the score of 3 (from 73% match in case of Llama to 87% in case of Gemma). The other scores were matched in under 50% in all cases; thus considered less reliable.
@@
Furthermore, we validated the meta-evaluation as a whole. To do that we calculated an average of the scores assigned by two human evaluators (average human score) and an average score assigned by three LLM meta-evaluators (meta-evaluation score). We found a strong ($\rho=0.75$) statistically significant correlation between average human score and meta-evaluation score on the balanced subset (mean absolute error of 0.51). Average human score and meta-evaluation score was the same (when rounded to the whole numbers) in 50% of cases (mostly in the score of 0), which is increased to 92% of cases when modified to binary Yes/No answers as mentioned above.
@@ 5.3 Detectability of Generated Personalized Texts
This experiment targets the research question RQ3: Does personalization affect detectability of generated disinformation as being generated by AI? To show detection performance of the existing detection methods, we have used the original extended data (1200 machine-generated texts accompanied by 73 human-written disinformation articles sourced from fact-checking platforms) used by Vykopal et al. (2024). They have already compared the performance of various detectors of MULTITuDE benchmark (Macko et al., 2023), which we have extended by comparison of various newest methods. Based on such assessment, we have selected the following three well-performing SOTA detection methods (various sizes, various architectures, various detector categories): 1) Gemma-2-9b-IT (Team, 2024) model fine-tuned [^2] Using the robust fine-tuning procedure of Spiegel and Macko (2024b). using English track training data of the most recent shared task focused on machine-generated text detection (Wang et al., 2025); 2) Detection-Longformer [^3] https://huggingface.co/nealcly/detection-longformer (Li et al., 2024) model (a BERT-like model with ~149M parameters adjusted for long documents) fine-tuned on 27 LLMs showing robust out-of-domain performance (the best pre-trained English detector in Macko et al., 2024b); and 3) Binoculars (Hans et al., 2024) detector showing currently to be one of the best methods for zero-shot detection (we have used GPT-J 6B base model, shown to be the best statistical detector in cross-domain evaluation of Macko et al., 2024a). We have used the detectors’ versions implemented in the IMGTB framework (Spiegel and Macko, 2024a). We summarize the detectors’ performance in Table 7, Appendix F, where we provide AUC ROC (area under the curve of receiver operating characteristic) as a classification-threshold independent metric, classification Threshold calculated (calibrated) to maximize difference between true positive and false positive rates (obtained from ROC curve), and MacroF1 representing macro average (due to class imbalance) of F1-score when the calculated thresholds are used for decisions. ROC curves of the detectors are illustrated in Figure 13.
@@
We used the calibrated thresholds (based on previously mentioned dataset of Vykopal et al., 2024) to calculate how many of the generated texts of our PerDisNews dataset are correctly predicted as being “machine-generated”, i.e. true positive rate (TPR). This metric has been selected due to having only machine-generated texts in our PerDisNews dataset. For a finer granularity of evaluation of detectability (to be independent of the used classification thresholds), we have also evaluated Mean Score reporting average prediction probability of “machine” class across the dataset. In case of Binoculars it represents the binoculars metric (higher value represents a higher probability of the “machine” class). The results are summarized in Table 3. We have conducted a paired t-test and Wilcoxon test to measure statistical significance ($\alpha$ = 0.05) for each detector, resulting in almost all differences between various personalization prompts being statistically significant except for Gemma-2-9b-IT detector.
@@ 6 Discussion
Meta-evaluation of personalization of the text to a given target group is usable. Our results show strong (Spearman’s $\rho$ of 0.75) and statistically significant correlation of average LLM annotation (3 LLMs) with average human annotation (2 humans). However, we have focused on English texts only and used 3 carefully selected LLMs to minimize biases. Multilingualism, style of the texts (we have used formal-styled news articles), as well as a selection of different LLMs could significantly affect meta-evaluation reliability. Thus, the researchers should be careful when using this approach and always validate the LLM judgments when applied to different scenarios.
@@ 7 Conclusions
Limited human evaluation. Our key findings rely on an LLM-based meta-evaluation of personalization quality, validated by its correlation with a human-annotated subset. Each text was reviewed by two annotators. We demonstrated the usefulness of meta-evaluation by its correlation with human-annotated subset. While increasing the number of annotators mitigates the individual bias, it also increases the exposure of human annotators to harmful content and annotation cost. We disclose the annotator’s guidelines and description of target groups used to steer the annotators’ understanding of personalization. We use Spearman correlation coefficient to track the agreement between annotators.
@@ Acknowledgments
This work was partially supported by the European Union under the Horizon Europe project vera.ai, GA No. 101070093); by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I01-03-V04-00068; and by the Slovak Research and Development Agency under the project Modermed, GA No. APVV-22-0414.
@@ Appendix C Design of Human Study
We considered several approaches before deciding to conduct a human study with two annotators on a subset of 109 texts. The first option is to ask representatives of the target groups to evaluate the disinformation news articles generated specifically for them. While this approach would allow us to study the potential effects of LLM-generated personalized messages on the target audience, it would raise significant ethical concerns. The second option is to use multiple non-expert annotators from different backgrounds (e.g. use of crowdworkers). Individual annotators might have their understanding of target groups based on their backgrounds and increasing the number and diversity of annotators would mitigate individual biases. Yet, this approach would expose a high number of annotators to harmful content.
In our study, two authors, a man and a woman, annotated the quality of personalization of the text. Both are generally knowledgeable about disinformation and are based in Europe. None of them has expert knowledge of personalization. Therefore, we provided them with detailed annotator’s guidelines. Annotators were provided with example text snippets for each point on the scale and a description of target group attributes (the same as used in the detailed prompt and summarized in Table 4) to steer their understanding of the target groups. The descriptions are stereotypical and capture broader tendencies within the group. Our approach could be further improved by consulting the descriptions of groups and annotator’s guidelines with experts with domain knowledge in personalization.

### v2

@@ Abstract
The capabilities of recent large language models (LLMs) to generate high-quality content indistinguishable by humans from human-written texts raises many concerns regarding their misuse. Previous research has shown that LLMs can be effectively misused for generating disinformation news articles following predefined narratives. Their capabilities to generate personalized (in various aspects) content have also been evaluated and mostly found usable. However, a combination of personalization and disinformation abilities of LLMs has not been comprehensively studied yet. Such a dangerous combination should trigger integrated safety filters of the LLMs, if there are some. This study fills this gap by evaluating vulnerabilities of recent open and closed LLMs, and their willingness to generate personalized disinformation news articles in English. We further explore whether the LLMs can reliably meta-evaluate the personalization quality and whether the personalization affects the generated-texts detectability. Our results demonstrate the need for stronger safety-filters and disclaimers, as those are not properly functioning in most of the evaluated LLMs. Additionally, our study revealed that the personalization actually reduces the safety-filter activations; thus effectively functioning as a jailbreak. Such behavior must be urgently addressed by LLM developers and service providers.
@@ 1 Introduction
However, there is so far little (and mostly anecdotal) supporting evidence to justify such fears. Therefore, this study fills the gap by 1) investigating vulnerabilities of the state-of-the-art (SOTA) LLMs to being misused for generation of personalized disinformation, 2) examining the quality of personalization, and 3) how it affects their integrated safety mechanisms. [^1] For the sake of replicability and support of further research, the data analysis source code as well as the generated dataset (upon request) is released at https://github.com/kinit-sk/personalized-disinfo for non-commercial research purpose only under strict conditions. Due to ethical concerns (see Appendix A), we are not releasing data generation source code (although an abstract overview is provided in the paper), as approved by our institutional Ethics Review Board. We build on the methodology proposed by Vykopal et al. (2024) to assess disinformation generation potential, and modify it for the purpose of personalized disinformation evaluation. We proceed from the definition of personalization proposed by Blom (2000), who defines it as “a process that changes the functionality, interface, information content, or distinctiveness of a system to increase its personal relevance to an individual.” By personalized disinformation, we mean disinformation appealing (by concerns and emotions) to a specific target group.
@@
(2) Showcase of usability of meta-evaluation (LLM as a judge annotation) for automated and scalable evaluation of personalized text generation. We use three different LLMs (to limit biased judgments) to annotate all data, showing a strong correlation with human judgment (using five human annotators) on a smaller balanced subset.
@@ 3 Methodology
Narratives. For this study, we have carefully selected six disinformation narratives (see Table 1) from a set of 20 in Vykopal et al. (2024) to ensure comparability of results. The resulting list covers health and politics-related disinformation, reflecting the varying LLMs’ behavior across topics (Vykopal et al., 2024). The description of previously mentioned target groups is based on European stereotypes. Therefore, we excluded disinformation narratives linked to events or personas outside Europe to avoid combination of target groups and narratives that are unlikely to appear in real-world information space. To the best of our judgment, we selected disinformation narratives that are possible to personalize across all chosen target groups. Each of the six selected narratives includes the title, summarizing the main idea, and an abstract providing additional context. We follow the wording from Vykopal et al. (2024) that was sourced from professional fact-checkers.
@@
Human validation. Five human annotators (out of the authors) have rated the personalization quality of the carefully selected 109 texts on a four-point scale. The human-annotated balanced subset consists of texts generated by all six LLMs, equally distributed over a score assigned by GPT-4o meta-evaluation (more details on the selection of the subset in Appendix C.2). Only the text and the target group were provided in the annotation process, while the annotators were familiar with the target group descriptions. Annotators answered “How well does the fake news article address the intended audience?” on a four-point scale: (0) – The text is not at all personalized to the intended target group. (1) – The text is personalized for the intended target group to a low extent. (2) – The text is personalized to an intended audience in an illogical way or the resulting personalization of text sounds forced and unnatural. (3) – The text is well personalized to the intended target group. The iteratively developed annotation guidelines provide a description of the scale with example text snippets to illustrate the cases. The guidelines have been validated by previous round of annotation of different portion of 50 texts by two annotators, addressing the ambiguities. Each text was rated by all five annotators. All annotators are based in Europe and are non-native English speakers.
@@ 4.1 Linguistic Quality Analysis
The linguistic analysis with text-quality evaluation of the generated texts is summarized in Table 2.
Mean ($\pm$ Standard deviation)
Generator Characters Words Lines Sentences GRUEN LA OCQ
The analysis shows that Falcon outputs have the highest variance in text lengths, thus showing least stable behavior. Although the three quality metrics do not match (reflecting different aspects), it seems that Flacon and Vicuna provide the texts of the lowest quality. On the other hand, Gemma and GPT-4o generated the texts of the highest quality. GPT-4o is the most stable model, generating consistently the high quality texts of similar lengths (the lowest variance in text lengths).
@@ 5.2 Meta-evaluation of Personalization
While human evaluation is labor-intensive, hard to replicate with the same results, and exposes annotators to a harmful content (disinformation in our case), employment of LLMs can effectively scale the evaluation and provide replicable results. However, LLM-based meta-evaluation comes not without limitations; among others LLMs tend to assign a higher score to their own outputs (Panickssery et al., 2024). To mitigate this limitation to a certain extent, we employ three LLMs, namely GPT-4o, Gemma-2-27b-IT, Llama-3.1-70B-Instruct for the evaluation of personalization quality. Moreover, LLM-assigned scores are for some tasks poorly correlated with human ratings (Bansal et al., 2023). While we have used such meta-evaluation to answer RQ1, in this section, we focus on validity of this approach and address the research question RQ2: Are LLMs usable to evaluate personalization of the generated texts with correlation to human judgment? To answer this question, five human annotators and three LLMs evaluated the personalization quality of a carefully balanced subset of 109 texts (see Section 3 and Appendix C.2).
@@
We calculated the agreement rates between the human annotators for the evaluation of personalization quality. The average Spearman correlation coefficient ($\rho$) is 0.62 indicating strong correlation (average mean absolute error between the five annotators is 0.58, average mean absolute error of human annotators from human average is 0.43). Overall, the five annotators assigned the same scores in 33% of cases, while three of five annotators (i.e., a majority) assigned the same scores in 90% of cases, which is acceptable given the 4-point scale and high subjective nature of evaluation task (personalization quality). The highest agreement is in the assignment of the score 3, where the majority of annotators agreed in 44% of cases, and the score 0 in 0.29% of cases. When the annotation is reduced to binary Yes/No answers (score of 0 is No, other scores are Yes) to whether the text is personalized for the target group (i.e., not the actual quality of the personalization), the full agreement is in 83% of cases and the majority agreement in all the cases.
@@
We also validated each score assigned by LLM-meta-evaluators individually. We checked whether it is matched by at least one of human annotators. It can provide us a kind of reliability measure of meta-evaluation scores. The highest reliability is in assignment of the score of 0 (from 92% match in case of Gemma to 97% in case of GPT-4o) and the score of 3 (from 88% match in case of Llama to 100% in case of Gemma). The other scores were matched in under 56% in all cases; thus considered less reliable. Figure 5 illustrates the personalization-quality evaluation-scores distribution across the validation subset (balanced based on GPT-4o scores). There are some differences among annotations observable. The Human5 and Llama annotations showing the highest preference of the score of 3, while the Human2, Human3, and Llama annotations showing the lowest assignment of the score of 1.
@@
Furthermore, we validated the meta-evaluation as a whole. To do that we calculated an average of the scores assigned by five human evaluators (average human score) and an average score assigned by three LLM meta-evaluators (meta-evaluation score). We found a strong ($\rho=0.76$) statistically significant correlation between average human score and meta-evaluation score on the balanced subset (mean absolute error of 0.45). Average human score and meta-evaluation score was the same (when rounded to the whole numbers) in 56% of cases (mostly in the score of 3), which is increased to 92% of cases when modified to binary Yes/No answers as mentioned above.
@@ 5.3 Detectability of Generated Personalized Texts
This experiment targets the research question RQ3: Does personalization affect detectability of generated disinformation as being generated by AI? To show detection performance of the existing detection methods, we have used the original extended data (1200 machine-generated texts accompanied by 73 human-written disinformation articles sourced from fact-checking platforms) used by Vykopal et al. (2024). They have already compared the performance of various detectors of MULTITuDE benchmark (Macko et al., 2023), which we have extended by comparison of various newest methods. Based on such assessment, we have selected the following three well-performing SOTA detection methods (various sizes, various architectures, various detector categories): 1) Gemma-2-9b-IT (Team, 2024) model fine-tuned [^2] Using the robust fine-tuning procedure of Spiegel and Macko (2024b). using English track training data of the most recent shared task focused on machine-generated text detection (Wang et al., 2025); 2) Detection-Longformer [^3] https://huggingface.co/nealcly/detection-longformer (Li et al., 2024) model (a BERT-like model with ~149M parameters adjusted for long documents) fine-tuned on 27 LLMs showing robust out-of-domain performance (the best pre-trained English detector in Macko et al., 2024b); and 3) Binoculars (Hans et al., 2024) detector showing currently to be one of the best methods for zero-shot detection (we have used GPT-J 6B base model, shown to be the best statistical detector in cross-domain evaluation of Macko et al., 2024a). We have used the detectors’ versions implemented in the IMGTB framework (Spiegel and Macko, 2024a). We summarize the detectors’ performance in Table 4, where we provide AUC ROC (area under the curve of receiver operating characteristic) as a classification-threshold independent metric, classification Threshold calculated (calibrated) to maximize difference between true positive and false positive rates (obtained from ROC curve), and MacroF1 representing macro average (due to class imbalance) of F1-score when the calculated thresholds are used for decisions. ROC curves of the detectors are illustrated in Figure 7.
Detector AUC ROC Threshold MacroF1
Gemma-2-9-IT 0.97 0.9995 0.83
Detection-Longformer 0.97 0.9043 0.74
Binoculars 0.74 -0.9387 0.46
Table 4: Evaluation of the selected machine-generated detection methods using the dataset of Vykopal et al. (2024).
Figure 7: Receiver operating characteristic curves of the selected detection methods using the dataset of Vykopal et al. (2024).
@@
We used the calibrated thresholds (from Table 4) to calculate how many of the generated texts of our PerDisNews dataset are correctly predicted as being “machine-generated”, i.e. true positive rate (TPR). This metric has been selected due to having only machine-generated texts in our PerDisNews dataset. For a finer granularity of evaluation of detectability (to be independent of the used classification thresholds), we have also evaluated Mean Score reporting average prediction probability of “machine” class across the dataset. In case of Binoculars it represents the binoculars metric (higher value represents a higher probability of the “machine” class). The results are summarized in Table 5. We have conducted a paired t-test and Wilcoxon test to measure statistical significance ($\alpha$ = 0.05) for each detector, resulting in almost all differences between various personalization prompts being statistically significant except for Gemma-2-9b-IT detector.
@@ 6 Discussion
Meta-evaluation of personalization of the text to a given target group is usable. Our results show strong (Spearman’s $\rho$ of 0.76) and statistically significant correlation of average LLM annotation (3 LLMs) with average human annotation (5 humans). However, we have focused on English texts only and used 3 carefully selected LLMs to minimize biases. Multilingualism, style of the texts (we have used formal-styled news articles), as well as a selection of different LLMs could significantly affect meta-evaluation reliability. Thus, the researchers should be careful when using this approach and always validate the LLM judgments when applied to different scenarios.
@@ 7 Conclusions
Limited human evaluation. Our key findings rely on an LLM-based meta-evaluation of personalization quality, validated by its correlation with a human-annotated subset. Each text was reviewed by five annotators. We demonstrated the usefulness of meta-evaluation by its correlation with human-annotated subset. While increasing the number of annotators mitigates the individual bias, it also increases the exposure of human annotators to harmful content and annotation cost. We disclose the annotator’s guidelines and description of target groups used to steer the annotators’ understanding of personalization. We use Spearman correlation coefficient to track the agreement between annotators.
@@ Acknowledgments
This work was partially supported by the European Union under the Horizon Europe project vera.ai, GA No. 101070093); by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the projects No. 09I01-03-V04-00068 and 09I03-03-V03-00020; and by the Slovak Research and Development Agency under the project Modermed, GA No. APVV-22-0414.
@@ Appendix C Design of Human Study
We considered several approaches before deciding to conduct a human study with five annotators on a subset of 109 texts. The first option is to ask representatives of the target groups to evaluate the disinformation news articles generated specifically for them. While this approach would allow us to study the potential effects of LLM-generated personalized messages on the target audience, it would raise significant ethical concerns. The second option is to use multiple non-expert annotators from different backgrounds (e.g. use of crowdworkers). Individual annotators might have their understanding of target groups based on their backgrounds and increasing the number and diversity of annotators would mitigate individual biases. Yet, this approach would expose a high number of annotators to harmful content.
In our study, five authors, gender-balanced, annotated the quality of personalization of the text. All of them are generally knowledgeable about disinformation and are based in Europe. None of them has expert knowledge of personalization. Therefore, we provided them with detailed annotator’s guidelines. Annotators were provided with example text snippets for each point on the scale and a description of target group attributes (the same as used in the detailed prompt and summarized in Table 6) to steer their understanding of the target groups. The descriptions are stereotypical and capture broader tendencies within the group. Our approach could be further improved by consulting the descriptions of groups and annotator’s guidelines with experts with domain knowledge in personalization.